OpenStudy (anonymous):

Solve the equation. 4p^2/5 - 6p^1/5 - 4 = 0 p= (smaller) p= (larger value)

OpenStudy (anonymous):

this is a quadratic equation in \[p^{\frac{1}{5}}\]

OpenStudy (anonymous):

solve \[4x^2-6x-4=0\] then replace \[x\] by \[p^{\frac{1}{5}}\]

OpenStudy (anonymous):

\[4x^2-6x-4=0\] \[2x^2-3x-2=0\] \[(x-2)(2x+1)=0\] \[x=2,x=-\frac{1}{2}\]

OpenStudy (anonymous):

\[p^{\frac{1}{5}}=2,p=2^5=32\]

OpenStudy (anonymous):

\[p^{\frac{1}{5}}=-\frac{1}{2},p=-\frac{1}{32}\]\]